Katy:

A circle with radii 12 mm is divided into 20 sectors of equal area. Which is the area of one sector to the nearest tenth? 1.9 square mm 3.8 square mm 22.6 square mm 90.5 square mm

Hero:

@Katy, we know the radius of the circle correct?

Hero:

12mm, yes. So this means we can calculate the area of the circle, correct?

Hero:

And you know how to calculate the area, right?

Hero:

Okay, so what is the area of the circle?

Hero:

If you leave off \(\pi\) then you're leaving off an important part of the area. \(144\pi\) is the area.

Hero:

So now, how do you suppose we might calculate the area of one sector of the circle now that we have the area of the whole circle?

Hero:

\(144 \times 3.14 \approx 144\pi\)

Hero:

Mind answering my question now?

Hero:

You have to multiply 144 by 3.14 to get a good approximation of \(\pi\)

Hero:

I'll ask this question again: How do you suppose we might calculate the area of one sector of the circle now that we have the area of the whole circle?

Hero:

Where's your calculator?

Hero:

@Ultrilliam do you see what this user just said?

Hero:

Which Operating System does your computer use?

Hero:

Most Operating Systems come with calculators pre-installed. All you would have to do is add it to your taskbar so that you can access it quickly

jorellcross:

i like but not that kind

jorellcross:

math

Hero:

There was only one step left, but @Katy doesn't want to finish.

Hero:

@Katy once you've found the area of the circle, the next step is to divide the circle into 20 sectors to find the area of one sector.